OpenStudy (anonymous):

Are f(x)=(2)/(3)x+6 and g(x)=(3)/(2)x-9 inverse of each other.

OpenStudy (anonymous):

no they aren't

OpenStudy (anonymous):

find f-1(x) and it wont equal g(x)

OpenStudy (anonymous):

f(x)=2/3x+6 An inverse of a function involves an x,y-interchange; that is, flip the x and y around and solve for y again. Thus, f(x) = y = 2 / (3x+6) x = 2/(3y+6) x(3y+6) = 2 3xy + 6x = 2 3xy = 2-6x y = (2-6x)/3x Since this doesn't match g(x), they are NOT inverses of one another. OR... do you mean f(x)= (2/3)x+6 = y x = (2/3)y + 6 x-6 = (2/3)y (x-6)(3/2) = y (3/2)x - 9 = y In this case, they ARE inverses of one another. I'm not sure what you really meant so I gave you to scenarios for f[x].

OpenStudy (anonymous):

This is how the question reads...Determine whether f(x) and g(x) are inverses of each other. f(x)=(2)/(3)x+6 g(x)=(3)/(2)x-9

OpenStudy (anonymous):

appreciate it so much, was just clarifiying the question:) Is the answer yes?

OpenStudy (anonymous):

Yes, :)

OpenStudy (anonymous):

is there something you don't understand ? ;o

OpenStudy (anonymous):

Yes, Alot lol but I'm learning as I go. It helps for me to figure HOW it come to that answer.